OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 3 + 3 + 9 + ... + 81

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

it is an A.P. a=-9 d=-3-(-9)=-3+9=6 find n by the the formula \[l=a+(n-1)d\] can you find n l=81

OpenStudy (anonymous):

are you here?

OpenStudy (anonymous):

these are answer options find n

OpenStudy (anonymous):

i did A @surjithayer

OpenStudy (anonymous):

i want to know how? can you show me ?

OpenStudy (anonymous):

@surjithayer i guessed im sry i suck at math

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

81=-9+(n-1)*6 81+9=(n-1)6 \[\frac{ 90 }{ 6 }=n-1\] n-1=15 n=15+1=16

OpenStudy (anonymous):

you got it.

OpenStudy (anonymous):

but on all of them n=0

OpenStudy (anonymous):

wait is it d?

OpenStudy (anonymous):

that means there are 16 terms for n=0,1,2,...,15

OpenStudy (anonymous):

here for A ,n=0 a1=-9 a2=-9+6*1=-3 a3=-9+6*2=3 ..... a16=-9+6*15=81

OpenStudy (anonymous):

I said A in the first place @surjithayer

OpenStudy (anonymous):

A is correct. as i have shown.

OpenStudy (dan815):

would you like to see how sum of arithmetic series formula comes

OpenStudy (anonymous):

could u show me on this one? Write the sum using summation notation, assuming the suggested pattern continues. 5 - 15 + 45 - 135 + ...

OpenStudy (anonymous):

it is a G.P. can you find r?

OpenStudy (anonymous):

what is gp, i need it in summation notation

OpenStudy (anonymous):

gp means geometric progression

OpenStudy (anonymous):

for that you have to find r(common ratio)

OpenStudy (anonymous):

\[r=\frac{ t _{n} }{ t _{n-1} }\] to be precise 2nd term/1st term

OpenStudy (anonymous):

@surjithayer what one did u get as the answer

OpenStudy (anonymous):

\[r=\frac{ 45 }{ -15 }=?\]

OpenStudy (anonymous):

\[r=\frac{ -15 }{ 5 }=\frac{ 45 }{ -15 }=?\]

OpenStudy (anonymous):

@surjithayer its b right?

OpenStudy (anonymous):

r=-3 B is correct.